# On the Interplay between Human Label Variation and Model Fairness

### v1

@@ Abstract
The impact of human label variation (HLV) on model fairness is an unexplored topic. This paper examines the interplay by comparing training on majority-vote labels with a range of HLV methods. Our experiments show that without explicit debiasing, HLV training methods have a positive impact on fairness.111Code and supplementary materials: https://github.com/kmkurn/hlv-fairness-interplay
@@ 1 Introduction
Human label variation (HLV; Plank, 2022) has been shown to improve performance through better generalisation (Peterson et al., 2019; Uma et al., 2021b; Kurniawan et al., 2025). However, the impact of HLV on model fairness remains unexplored. It is well established that there is a trade-off between performance and fairness (Han et al., 2022; Shen et al., 2022), so one might conjecture that HLV would similarly negatively impact fairness (since HLV improves performance). We explore this interplay between HLV and model fairness in this paper, in what we believe to be the first systematic analysis of this interaction. We compare training on majority-vote labels with four HLV training methods from prior work (Kurniawan et al., 2025), and evaluate on two tasks: offensiveness classification with SBIC (Sap et al., 2020) and legal area classification with TAG, an in-house legal dataset.
We find that across SBIC and TAG, training with HLV methods improves overall performance without harming fairness. Over TAG, training with HLV even improves fairness. Since these findings are only from a single set of class and group weights (configuration hereafter), we examine many possible configurations to draw more robust conclusions. For both datasets, HLV does not reduce model fairness for most configurations. Interestingly, HLV can improve model fairness for a substantial number of configurations.
We then investigate why HLV improves fairness by introducing temperature scaling to an HLV training method to control the weighting of minority annotations. Our TAG experiments suggest fairness gains indeed come from minority annotations.
Our findings suggest that training with HLV improves performance without compromising fairness. That said, our findings also underscore the effect of class and group weightings in measuring fairness in the HLV context. Therefore, it is important to understand what fairness means for each application before incorporating HLV.
@@ 3.2 Evaluation
To evaluate fairness, we first compute a fairness score (defined later) for each class and group . Next, following Han et al. (2023), we aggregate over groups then classes using weighted generalised mean to obtain the overall fairness score. The group-wise aggregation is defined as
@@
For TAG results in Section˜4.1, the class and group aggregation weights are provided by the legal organisation who co-developed the dataset (Section˜3.1). For SBIC, as we have no a priori preference we weight each class and group equally. For both datasets, we set for both aggregation, resulting in a standard weighted average. In Section˜4.2, we experiment with multiple randomly-sampled configurations for more robust conclusions.101010Details are given in Appendix E.
@@ 3.3 Approaches
For HLV training methods, we use the four most successful methods identified by Kurniawan et al. (2025): repeated labelling (ReL), soft labelling (SL), minimising Jensen-Shannon divergence (JSD), and maximising soft micro F1 score (SmF1).131313See Appendix F for more details. For non-HLV baselines, we finetune them on majority-voted classes (MV).141414Further training details are reported in Appendix G.
@@ 4.1 Single Fairness Configuration
We report the overall performance and fairness of each method in Table˜1. For SBIC, the table shows that HLV training methods consistently outperform MV for overall performance across the two models. These improvements are statistically significant (with bootstrap tests) for 6 out of 8 model–method pairs.151515Exceptions: SL and JSD with RoBERTa As for fairness, the HLV training methods can improve or reduce fairness, but the differences are not statistically significant for 7 out of 8 model–method pairs.161616Exception: SmF1 with RoBERTa
Looking at the TAG results, for both overall performance and fairness, the HLV training methods consistently outperform MV across both models, with statistically significant improvements for 6 out of 8 model–method pairs.171717Exception: SmF1 Taken together, the results suggest that HLV training improves overall performance without sacrificing fairness, and in some cases (e.g., TAG), can even improve fairness.
@@ 4.2 Multiple Fairness Configurations
The results in Section˜4.1 were obtained with one configuration, i.e. fixed values for and in Equation˜1 for both group- and class-wise aggregation. Here, we test multiple randomly-sampled configurations to draw more robust conclusions.181818See Appendix E for more details. This experimental setup is especially important for SBIC: the single configuration used in Section˜4.1 is divorced from any intended application of the model.
For each randomly-sampled configuration,191919Note that performance is constant regardless of configurations as they configure only the fairness score computation. we perform a two-tailed bootstrap test across 3 runs to check if the mean fairness score difference between each HLV training method and MV is statistically significant (). If so, we record if the mean fairness score of MV is higher (i.e. fairer). Otherwise, we deem the HLV training method as fair as MV. We report the fraction of configurations where MV is found not fairer than each HLV method on SBIC in Figure˜1. TAG follows this trend.202020The comparisons are shown in Appendix H.
@@ 5.2 Minority Annotations
Figure˜3 shows when is small, performance and fairness are low. As increases, performance gradually climbs but peaks at , then declines as grows further. In contrast, fairness improves as grows, then stabilises after peaking at . These findings suggest minority annotations contain useful signals for fairness.232323We offer a possible explanation in Appendix J. Unsurprisingly, results show performance peaks at because the scaling is only applied to the train set, i.e. test ground truths are computed normally ().
@@ 6 Conclusions
While HLV has been shown to improve generalisation, this is the first work to study the interplay between HLV and fairness. We first showed empirically that training with HLV does not harm model fairness, and indeed, it can improve fairness. We then showed that this observation is robust to various fairness configurations. However, when the lowest-scoring groups or classes in fairness were prioritised, training with HLV provided no fairness benefits compared to training on majority-voted classes. Lastly, we confirmed that the fairness improvements with HLV training come from the minority annotations. Our findings highlight the performance and fairness benefits of training with HLV, but we recommend NLP practitioners to consider what fairness means in their application before incorporating HLV.
@@
Because the TAG dataset contains real-world confidential legal help requests, we cannot distribute it as it used in the study. However, we still include it because of its importance to our analysis in Section˜5.2, and we believe that the insight obtained from the dataset in the study is still useful. Although we cannot distribute the dataset, we can distribute ground-truth and predicted judgement distributions of the dev and the test sets.242424https://github.com/kmkurn/hlv-fairness-interplay This will allow other researchers to replicate our reported numbers and build on our work to some extent.
We evaluate using soft evaluation metrics in all experiments. One could argue that this setup favours HLV training, so the comparisons with MV are unfair. However, we note that previous HLV work has found that MV outperforms many HLV training methods in this setup (Kurniawan et al., 2025), suggesting that the argument doesn’t hold in practice.
@@ Acknowledgments
This research is supported by the Australian Research Council Linkage Project LP210200917252525https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government. This research is done in collaboration with Justice Connect, an Australian public benevolent institution.262626As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project. This research is supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.
@@ Appendix G Training Details
We implement all methods using FlairNLP (Akbik et al., 2019).272727https://flairnlp.github.io For RoBERTa, we tune the learning rate and the batch size using random search, optimising only for overall performance (without fairness). Table˜3 reports the best values found. For OLMo, we use FlairNLP’s default hyperparameters282828Learning rate and batch size are and 32 respectively. for computational reasons. We use LoRA (Hu et al., 2022) to finetune OLMo efficiently. All models are finetuned for 10 epochs. On a single NVIDIA A100 GPU with 80GB of memory, the finetuning runs for:
@@
Appendix H More Multiple Fairness Configurations Results
@@
Appendix I More Fairness Configurations Analysis Results
@@
Appendix J Possible Explanation on Minority Annotations Improving Fairness on TAG

### v2

@@ Abstract
The impact of human label variation (HLV) on model fairness is an unexplored topic. This paper examines the interplay by comparing training on majority-vote labels with a range of HLV methods. Our experiments show that without explicit debiasing, HLV training methods have a positive impact on fairness under certain configurations.111Code and supplementary materials: https://github.com/kmkurn/hlv-fairness-interplay
@@ 1 Introduction
Human label variation (HLV; Plank, 2022) has been shown to improve performance through better generalisation (Peterson et al., 2019; Uma et al., 2021b; Kurniawan et al., 2025). However, the impact of HLV on model fairness remains unexplored. One might conjecture that HLV would improve fairness because it preserves minority views and avoids collapsing them into a single point like majority voting. However, it is well established that there is a trade-off between performance and fairness (Han et al., 2022; Shen et al., 2022), so one might also conjecture that HLV would similarly negatively impact fairness (since HLV improves performance). We explore this interplay between HLV and model fairness in this paper, in what we believe to be the first systematic analysis of this interaction. We compare training on majority-vote labels with four HLV training methods from prior work (Kurniawan et al., 2025), and evaluate on two tasks: offensiveness classification with SBIC (Sap et al., 2020) and legal area classification with TAG, an in-house legal dataset, developed with practising lawyers.
We find that across SBIC and TAG, training with HLV methods improves overall performance without harming fairness. Over TAG, training with HLV even improves fairness. Since these findings are only from a single set of class and group weights (configuration hereafter), we examine many possible configurations to draw more robust conclusions. For both datasets, HLV does not reduce model fairness for most configurations, and this trend is clearer when the configuration prioritises all groups/classes equally. Interestingly, HLV can improve model fairness for a substantial number of configurations.
We then investigate why HLV improves fairness by introducing temperature scaling to an HLV training method to control the weighting of minority annotations. Our TAG experiments suggest fairness gains do indeed come from minority annotations.
Our findings suggest that training with HLV improves performance without compromising fairness. Additionally, our findings underscore the effect of class and group weightings in measuring fairness in the HLV context. Therefore, it is important to understand what fairness means for each application before incorporating HLV.
@@ 3.2 Evaluation
We focus on a group-level parity definition of fairness in this study due to its prevalence in prior work (Shen et al., 2022; Han et al., 2023). To evaluate this, we first compute a fairness score (defined later) for each class and group . Next, following Han et al. (2023), we aggregate over groups then classes using weighted generalised mean to obtain the overall fairness score. The group-wise aggregation is defined as
@@
For TAG results in Section˜4.1, the class and group aggregation weights are provided by the legal organisation who co-developed the dataset (Section˜3.1). For SBIC, as we have no a priori preference we weight each class and group equally. For both datasets, we set for both aggregations, resulting in standard weighted average. In Section˜4.2, we experiment with multiple randomly-sampled configurations for more robust conclusions.101010Details are given in Appendix E.
@@ 3.3 Approaches
For HLV training methods, we use the four most successful methods identified by Kurniawan et al. (2025): repeated labelling (ReL), soft labelling (SL), minimising Jensen-Shannon divergence (JSD), and maximising soft micro F1 score (SmF1).131313See Appendix F for more details. For non-HLV baselines, we finetune them on majority-vote classes (MV).141414Further training details are reported in Appendix G.
@@ 4.1 Single Fairness Configuration
We report the overall performance and fairness of each method in Table˜1. For SBIC, the table shows that HLV training methods consistently outperform MV for overall performance across the two models. These improvements are statistically151515We report effect sizes as a measure of practical significance in Appendix H. significant (with bootstrap tests) for 6 out of 8 model–method pairs.161616Exceptions: SL and JSD with RoBERTa As for fairness, the HLV training methods can improve or reduce fairness, but the differences are not statistically significant for 7 out of 8 model–method pairs.171717Exception: SmF1 with RoBERTa
Looking at the TAG results, for both overall performance and fairness, the HLV training methods consistently outperform MV across both models, with statistically significant improvements for 6 out of 8 model–method pairs.181818Exception: SmF1 Taken together, the results suggest that HLV training improves overall performance without sacrificing fairness, and in some cases (e.g., TAG), can even improve fairness.
Discussion
Unsuprisingly, the different HLV methods attain different levels of fairness. However, these variations are consistent with their overall performance. For example, ReL, a strong method performance-wise, is also among the strongest in terms of fairness, while SmF1 is among the weakest. That SmF1 attains poor performance is not surprising because Kurniawan et al. (2025) showed that the method can result in degenerate predictions where all probability mass is concentrated on the majority class. Looking into class- and group-wise performance for each method, we find that all noticeable differences seem to align well with overall performance trends.
@@ 4.2 Multiple Fairness Configurations
The results in Section˜4.1 were obtained with one configuration, i.e. fixed values for and in Equation˜1 for both group- and class-wise aggregation. Here, we test multiple randomly-sampled configurations to draw more robust conclusions.191919See Appendix E for more details. This experimental setup is especially important for SBIC: the single configuration used in Section˜4.1 is divorced from any intended application of the model.
For each randomly-sampled configuration,202020Note that performance is constant regardless of configurations as they configure only the fairness score computation. we perform a two-tailed bootstrap test across 3 runs to check if the mean fairness score difference between each HLV training method and MV is statistically significant (). If so, we record if the mean fairness score of MV is higher (i.e. fairer). Otherwise, we deem the HLV training method as fair as MV. Note that overall performance is unaffected by fairness configurations. Therefore, we compare only the fairness scores.
We report the fraction of configurations where MV is found not to be fairer than each HLV method on SBIC in Figure˜1. The result on TAG follows the same trend, which we show in Appendix˜I.
@@ 5.2 Minority Annotations
We report how fairness changes as grows in Figure˜3. The figure shows that when is small, fairness is low. As increases, fairness improves and stabilises after peaking at . These findings suggest minority annotations contain useful signals for fairness. We offer a possible explanation for this in Appendix˜K and leave the full investigation on this direction for future work.
Unsurprisingly, the figure also shows that performance peaks at . This is because the scaling is only applied to the train set, i.e. test ground truths are computed normally ().
@@ 6 Conclusions
While HLV has been shown to improve generalisation, this is the first work to study the interplay between HLV and fairness. We first showed empirically that training with HLV does not harm model fairness, and indeed, it can improve fairness. We then showed that this observation is robust to various fairness configurations. However, when the lowest-scoring groups or classes in fairness were prioritised, training with HLV provided no fairness benefits compared to training on majority-vote classes. Lastly, we confirmed that the fairness improvements with HLV training come from the minority annotations.
Our findings highlight the performance and fairness benefits of training with HLV under certain configurations. Therefore, we recommend NLP practitioners to: (1) consider what fairness means in their application, (2) define it in terms of the fairness evaluation configuration described in Section˜3.2, (3) incorporate HLV if the configuration prioritises all groups/classes roughly equally, and (4) evaluate fairness with the configuration to confirm that HLV does not harm fairness compared to majority voting.
@@
Because the TAG dataset contains real-world confidential legal help requests, we cannot distribute it as it used in the study. However, we still include it because of its importance to our analysis in Section˜5.2, and we believe that the insight obtained from the dataset in the study is still useful. Although we cannot distribute the dataset, we can distribute ground-truth and predicted judgement distributions of the dev and the test sets.232323https://github.com/kmkurn/hlv-fairness-interplay This will allow other researchers to replicate our reported numbers and build on our work to some extent.
Prior work has evaluated fairness using synthetic, targeted check lists (Manerba and Tonelli, 2021). Therefore, a reasonable solution may be to follow a similar approach to address the data privacy issue. We agree in theory that it would allow us to not only measure fairness at a high level but also understand in what cases the models break down. However, our goal is to understand fairness and its relationship with HLV rather than address specific impacts of model bias. Furthermore, we foresee a few complications: (1) it is not always immediately clear what the biases might be as we need to know them in order to create the synthetic check lists); (2) in our case, the language of the TAG data is non-standard English, and so it is difficult to mimic that genre; and (3) we will need to distribute our TAG-trained models, which is restricted because of IP constraints and possible risks of leaking the sensitive TAG training data through the model. Therefore, this work does not take this approach.
We evaluate using soft evaluation metrics in all experiments. One could argue that this setup favours HLV training, so the comparisons with MV are unfair. However, we note that previous work has found that MV outperforms many HLV training methods in this setup (Kurniawan et al., 2025), suggesting that the argument does not hold in practice.
@@ Acknowledgments
This research is supported by the Australian Research Council Linkage Project LP210200917242424https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government. This research is done in collaboration with Justice Connect, an Australian public benevolent institution.252525As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project. This research is supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.
@@ Appendix G Training Details
We implement all methods using FlairNLP (Akbik et al., 2019).262626https://flairnlp.github.io For RoBERTa, we tune the learning rate and the batch size using random search, optimising only for overall performance (without fairness). Table˜3 reports the best values found. For OLMo, we use FlairNLP’s default hyperparameters272727Learning rate and batch size are and 32 respectively. for computational reasons. We use LoRA (Hu et al., 2022) to finetune OLMo efficiently. All models are finetuned for 10 epochs. On a single NVIDIA A100 GPU with 80GB of memory, the finetuning runs for:
@@
Appendix H Effect Sizes
[…]
| Model | Method | SBIC | TAG | ||||||
| Perf | Fair | Perf | Fair | ||||||
| low | high | low | high | low | high | low | high | ||
| OLMo | ReL | ||||||||
| SL | |||||||||
| JSD | |||||||||
| SmF1 | |||||||||
| RoBERTa | ReL | ||||||||
| SL | |||||||||
| JSD | |||||||||
| SmF1 | |||||||||
Appendix I More Multiple Fairness Configurations Results
@@
Appendix J More Fairness Configurations Analysis Results
@@
Appendix K Possible Explanation on Minority Annotations Improving Fairness on TAG